# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ 5.2.1 Maximum Independent Set Problem
Given a graph $G(\mathcal{N},\mathcal{E})$, an independent set is a subset of vertices where no two vertices are connected by an edge. A maximal independent set (MIS) is the largest independent set of $G$. Namely, we solve the optimization problem
$\mathcal{C}^{*}_{G=(\mathcal{N},\mathcal{E})}:=\min_{x\in\{0,1\}^{|\mathcal{N}|}}\left\{-\sum_{i\in\mathcal{N}}x_{i}:x_{i}+x_{j}\leq 1~{}\forall(i,j)\in\mathcal{E}\right\}.$ (Maximum Independent Set)
@@
The Hamiltonian $H$ is defined on constrained space $\mathcal{X}$ and therefore we need a constrained walk to explore the state space. For this purpose, we use the Glauber dynamics defined as
$P_{\lambda}(x,x^{\prime})=\begin{cases}0&\text{if }|x-x^{\prime}|>1\\ \frac{1}{n}\frac{\lambda}{\lambda+1}&\text{if }|x-x^{\prime}|=1\text{ and }x\subseteq x^{\prime}\\ \frac{1}{n}\frac{1}{\lambda+1}&\text{if }|x-x^{\prime}|=1\text{ and }x^{\prime}\subseteq x\\ 1-\sum_{x^{\prime\prime}\neq x}P_{\lambda}(x,x^{\prime\prime})&\text{if }x=x^{\prime}.\end{cases}$ (42)
This model is also referred as hard-core model in statistical physics and we’ll use the same terminology. Note that Glauber dynamics initialized at an independent set can only move between independent sets. Furthermore, it converges to its stationary distribution
@@
The parameter $\lambda$ is also called fugacity and one can recover the Gibbs form in 5.2 by setting $e^{\beta}=\lambda$. In principle, one can find the maximum independent set by setting $\lambda$ sufficiently high so that distribution $\pi_{\lambda}$ concentrates around the global minimum of $H$. Unfortunately, this approach results in exponential mixing time due to uniqueness/non-uniqueness phase transitions. Alternatively, one can draw exponentially many samples from $\pi_{\lambda}$ at $\lambda<\lambda_{c}$ by running the Glauber dynamics chain in polynomial time as Glauber dynamics mix efficiently below the critical threshold [CLV21]. As the current quantum techniques can only quadratically improve the run time of the first approach, we use the second apprach. More specifically, we consider Glauber chain $P_{\lambda}$ at $\lambda<\lambda_{c}$ so that we can prepare |$\ket{\pi_{\lambda}}$ efficiently. Next, we consider the short-path Hamiltonian $H_{b}:\mathcal{X}\mapsto\mathbb{R}$,
@@
where $D$ is the discriminant matrix as usual and $E^{\ast}=\min_{x}H(x)=-|x^{\ast}|$. In accordance with the generalized short path framework, the algorithm starts from |$\ket{\pi_{\lambda}}$ and jumps to the ground state of $H_{b}$ for $b>0$. We describe how we can prepare the block-encoding of $D$ for Glauber dynamics and also prepare $\pi_{\lambda}$ in Appendix D.
The following lemma establishes the condition on $\|H\|_{P}$ given in theorem 5.7.
@@
Let $P$ be a Glauber dynamics chain with fugacity parameter $\lambda$ on a graph $G(\mathcal{N},\mathcal{E})$. Then, $\|H\|_{P}=\mathcal{O}(1)$.
@@
As required by Theorem 5.7, we need to characterize the log-Sobolev constant of Glauber dynamics for hard-core model on a graphs with a degree upper bounded by $d$. We start with the following fact from [CLV21],
Fact 5.10.
Let $V$ be a set of size $n$ and µ be a distribution over $[q]^{V}$. If $\pi$ satisfies the approximate tensorization of entropy with constant $C_{1}$ and $\pi$ is $u$-marginally bounded, then the Glauber dynamics for $\pi$ satisfies the standard log-Sobolev inequality with constant $\omega=\frac{1-2u}{\log(1/u-1)}\frac{1}{C_{1}n}$ when $u<\frac{1}{2}$, or $\omega=\frac{1}{2C_{1}n}$ when $u=\frac{1}{2}$.
@@
Fix a configuration $x$ and consider a Markov chain $P_{v}$ that updates vertex $v$ according to marginal probability distribution $\pi_{v}=\pi(v|x_{V-\{v\}})$. Then, LS constant of this Markov chain $\rho_{v}$ is lower bounded by $\rho_{v}\geq\frac{1-2\pi_{v}^{\ast}}{\log(1/\pi_{v}^{\ast}-1)}$ when $\pi_{v}^{\ast}<\frac{1}{2}$ or $\rho_{v}=\frac{1}{2}$ when $\pi_{v}^{\ast}=\frac{1}{2}$ due to [DSC96, Theorem A.1]. By the definition of the log-Sobolev constant, we have
The transition matrix of Glauber dynamics can be written as $\frac{1}{n}\sum_{v\in V}\pi(v|\cdot)$. Therefore, the Dirichlet form for Glauber dynamics is
$\mathcal{D}(\sqrt{f},\sqrt{f})=\frac{1}{n}\sum_{v\in V}\mathrm{Var}_{\pi_{v}}[\sqrt{f}].$
Using the tensorization of entropy,
$\operatorname{Ent}_{\pi}[f]$ $\leq C_{1}\sum_{v\in V}\operatorname{Ent}_{\pi_{v}}[f]\leq C_{1}\sum_{v\in V}\frac{1}{\rho_{v}}\mathrm{Var}_{\pi_{v}}[\sqrt{f}]\leq C_{1}n\max_{v}\left({\frac{1}{\rho_{v}}}\right)\mathcal{D}(\sqrt{f},\sqrt{f}).$
Combining the marginal boundedness property of $\pi$ (i.e., $\pi_{v}\geq u$ for all $v\in V$), with the monotonicity of the function $\frac{1-2y}{\log(1/y-1)}$ for $y\in[0,1/2]$, it follows that $\omega\geq\frac{1-2u}{\log(1/u-1)}\frac{1}{C_{1}n}$ when $u<\frac{1}{2}$, and $\omega=\frac{1}{2C_{1}n}$ when $u=\frac{1}{2}$.
Theorem 5.11 (Entropy factorization, Theorem 2.9 in [CLV21]).
Let $d\geq 3$ be an integer and $b,\eta>0$ be reals. Suppose that $G=(\mathcal{N},\mathcal{E})$ is an n-vertex graph of maximum degree at most $d$ and $\mu$ is a totally connected Gibbs distribution of some spin system on $G$. If $\mu$ is both $u$-marginally bounded and $\eta$-spectrally independent and $n\geq\frac{24d}{u^{2}}(\frac{4\eta}{u^{2}}+1)$, then $\mu$ satisfies the approximate tensorization of entropy with constant
$C_{1}=\frac{18\log(1/u)}{u^{4}}\left(\frac{24d}{u^{2}}\right)^{\frac{4\eta}{u^{2}}+1}.$
Spectral independence and marginal boundedness properties for graphs with constant maximum degree are proven in [CLV23, CLV21] respectively. This shows that the log-Sobolev constant of a graph with bounded degree is $\Omega(n^{-1})$. Having showed that Glauber dynamics chain satisfies desired Log-Sobolev constant and $\|H\|_{P}$, we present the final run time.
Theorem 5.12.
Let $P$ be a Glauber dynamics chain on a graph $G(\mathcal{N},\mathcal{E})$ with a bounded maximum degree $d$ at fugacity parameter $\lambda<\lambda_{c}$ and stationary distribution $\pi$. Then, there exists a short-path algorithm that finds the maximum independent set in $G$ with running time
$\mathcal{O}^{\ast}\left([\pi(E^{\ast})^{-1}]^{\frac{1}{2}-c}\right),$
@@
It is evident that for sparse graphs ($d=\mathcal{O}(1)$), the size of the maximum independent set is $\Theta(n)$. Furthermore, $\log(1/\pi_{\lambda}(E^{\ast}))=\Theta(n)$ due to the hardness of finding the maximum independent sets. As log-Sobolev constant $\omega^{-1}=\mathcal{O}(n)$ for the Glauber dynamics on bounded degree graphs below the critical fugacity $\lambda_{c}$. Therefore, the conditions in Theorem 5.7 holds. Hence, we obtain a super-quadratic speedup over sampling from $\pi_{\lambda}$. ∎
Finally, we show that on a random regular graph, Markov Chain search using Glauber dynamics is better than brute force search.
@@
Let $\pi_{\lambda}$ be the stationary distribution of Glauber dynamics on a random $d$-regular graph $G\sim\mathcal{G}(n,d)$ with $d\geq d_{0}$. Choose $x^{\ast}$ to be a particular maximum independent set in $G$. Then, for $\lambda=\lambda_{c}$ we have
$\pi(\mathcal{I}^{\ast})\geq 2^{-\kappa n},$
with $\kappa=\frac{2\log(\lambda)\log d_{0}}{d_{0}}-\frac{1}{2d_{0}}-\frac{\log(1+\lambda)}{2}$.
@@
Now define $f(d)=\frac{2\log(\lambda)\log d}{d}-\frac{1}{2d}-\frac{\log(1+\lambda)}{2}$, where $\lambda=\frac{(d-1)^{(d-1)}}{(d-2)^{d}}$. Then, for $d>d_{0}$, we have $f(d)>-1$. ∎
@@
The parameter $\kappa$ from Proposition 5.13 is a decreasing function of $d_{0}$. Therefore, the performance of this algorithm improves as the degree of the graph increases.
5.2.2 Ising Model
Consider the 2-spin Ising model on a graph $G(\mathcal{N},\mathcal{E})$ defined via the Hamiltonian
$H(x)=\sum_{(i,j)\in\mathcal{E}}x_{i}x_{j}+\sum_{j}h_{j}x_{j},$
@@ 5.2.2 Ising Model
Assume $J_{ij}=1$ for all $(i,j)\in\mathcal{E}$ and $h=\mathbf{0}$. This model corresponds to Gibbs sampling with weights $\pi(x)\propto\exp(-\beta H(x))$, and a Gibbs sample can be prepared by using Glauber dynamics similar to the hardcore model. The only change is the transition probabilities, which can be computed by the marginal distribution $\pi(x_{i}^{t+1}|x^{t}_{\mathcal{N}\setminus\{i\}})$. For simplicity we consider the anti-ferromagnetic model, where having two neighboring sites have the same spin results in lower probability than having the same spin.
Assuming the underlying graph is sparse ($d=\mathcal{O}(1)$), then Glauber dynamics mixes in $\operatorname{poly}(n)$ time for $\beta<\frac{d-2}{d}=\beta_{c}$ [CLV21] at which the uniqueness/non-uniqueness phase transition occurs. We consider the following short path Hamiltonian
@@
where $P_{\beta}$ is the Glauber dynamics transition matrix for $\beta<\beta_{c}$. Similar to the setting of the MIS problem, a block-encoding of $D(P_{\beta})$ can be prepared efficiently.
Proposition 5.15.
The optimum energy of Ising Model Hamiltonian $H$ on a random regular graph $G(\mathcal{N},\mathcal{E})$ satisfies $|E^{\ast}|=\Theta(n)$.
@@
[…]
$|\textrm{BW}|=\frac{s+E_{\textrm{gs}}}{2}.$
Theorem 5.16.
Let $P$ be a Glauber dynamics chain on a random regular graph $G(\mathcal{N},\mathcal{E})$ with constant degree $d$ at inverse temperature parameter $\beta<\beta_{c}$ and stationary distribution $\pi_{\beta}$. Then there exists a short-path algorithm that finds the optimum of Ising model Hamiltonian on a random regular graph with running time
@@
As spectral independence and marginal boundedness for Ising model on sparse graphs are proven in [CLV21]. Therefore, Glauber dynamics for Ising model on a regular graph has $\omega^{-1}=\mathcal{O}(n)$. Furthermore, $\log(1/\pi_{\beta}(E^{\ast}))=\Theta(n)$ due to hardness of the problem. Similar to MIS problem, for all $x,x^{\prime}\in\{-1,1\}^{n}$ such that $P(x,x^{\prime})>0$, $\lvert H(x)-H(x^{\prime})\rvert\leq 2d=\mathcal{O}(1)$. By Proposition 5.15, $|E^{\ast}|=\mathcal{O}(n)$. Therefore, the conditions in Theorem 5.7 are satisfied. Hence, we have the super-quadratic sampling over sampling from $\pi_{\beta}$. ∎

### v2

@@ 5.2.2 Hardcore Model – Maximum Independent Set Problem
Given a graph $G=G(\mathcal{N},\mathcal{E})$, an independent set is a subset of vertices in $\mathcal{N}$ such that no two vertices are connected by an edge in $\mathcal{E}$. A maximal independent set (MIS) is the largest independent set of $G$. Namely, we solve the optimization problem
$\mathcal{C}^{*}_{G=(\mathcal{N},\mathcal{E})}:=\min_{x\in\{0,1\}^{|\mathcal{N}|}}\left\{-\sum_{i\in\mathcal{N}}x_{i}:x_{i}+x_{j}\leq 1~\forall(i,j)\in\mathcal{E}\right\}$ (Maximum Independent Set)
@@
The Hamiltonian $H$ is defined on the constrained space $\mathcal{X}$ and therefore we need a constrained walk to explore the state space. For this purpose, we use the Glauber dynamics defined as
[…]
This model is also referred as hardcore model in statistical physics and we’ll use the same terminology. Note that Glauber dynamics initialized at an independent set can only move between independent sets. Furthermore, it converges to its stationary distribution
@@
[…]
@@
[…]
The following lemma establishes a bound on $\|H\|_{P}$ for Glauber dynamics.
@@
Let $P$ be a Glauber dynamics chain on the Hardcore model with fugacity parameter $\lambda$ on a graph $G(\mathcal{N},\mathcal{E})$. Then, $\|H\|_{P}=\mathcal{O}(1)$.
@@
We have the following bound on the LS constant using the techniques from Section 5.2.1.
Corollary 5.16 (Log-Sobolev Constant For Constant-Fraction From Criticality).
Let $G=(\mathcal{N},\mathcal{E})$ be any graph with constant-bounded degree $d$, $n$ vertices, and $\lambda_{c}$ the corresponding tree-uniqueness threshold for the hardcore model. Under the conditions of Theorem 5.12, for any constant $\xi\in(0,1)$ and $\lambda\leq(1-\xi)\lambda_{c}$, the log-Sobolev constant of Glauber dynamics satisfies $\omega=\Omega(1/n)$.
@@
In [CLV23], as discussed in [CLV21, Proof of Theorem 1.1], it was shown that for any $\xi\in(0,1)$, $\lambda\leq(1-\xi)\lambda_{c}$ and that the spectral independence is $\rho=\mathcal{O}(1/\xi)$. Also, [CLV21, Proof of Theorem 1.1] showed that $u$ is lower bounded by a constant in $\xi$ and $\lambda$. Hence for graphs of constant maximum degree $d$, Theorem 5.12 and Fact 5.10 imply that the Glauber dynamics at fugacity $\lambda$ has a log-Sobolev constant that is $\Omega(1/n)$ for bounded-degree graphs. ∎
Having showed that Glauber dynamics chain satisfies desired log-Sobolev constant and bounded $\|H\|_{P}$, we present the final run time.
Theorem 5.17 (Constant Fraction From Criticality).
Let $G=(\mathcal{N},\mathcal{E})$ be any graph with a constant maximum degree $d$ and $\lambda_{c}$ the corresponding tree-unqiueness threshold for $d$. Let $P$ be a Glauber dynamics chain on a graph $G$ at fugacity parameter $\lambda\leq(1-\xi)\lambda_{c}$ for any constant $\xi\in(0,1)$ and stationary distribution $\pi_{\lambda}$. Then, there exists a short path algorithm that finds the maximum independent set in $G$ with running time
$\mathcal{O}^{\ast}\left([\pi_{\lambda}(E^{\ast})^{-1}]^{\frac{1}{2}-c}\right),$
@@
It is evident that for sparse graphs ($d=\mathcal{O}(1)$), the size of the maximum independent set is $\Theta(n)$. Since $d$ is constant in $n$, the problem still can not be solved in $\operatorname{poly}(n)$ time and requires $2^{\mathcal{O}(n)}$ samples for hard instances. Hence, $\log(1/\pi_{\lambda}(E^{\ast}))=\Theta(n)$ . The log-Sobolev constant from Corollary 5.16 satisfies $\omega^{-1}=\mathcal{O}(n)$ for the Glauber dynamics on bounded degree graphs for any constant-fraction below the critical fugacity $\lambda_{c}$. Also Lemma 5.15 implies a constant $P$-psuedo Lipschitz norm. Therefore, the conditions of Corollary 4.5 hold. Hence, we obtain a super-quadratic speedup over sampling from $\pi_{\lambda}$. ∎
The above result only applies for a constant fraction from the critical threshold $\lambda_{c}$, i.e. $\xi$ is constant. As mentioned earlier, Glauber dynamics is known to mix in polynomial time when $\xi=0$. However, it appears that the MLS and LS vanish at this threshold, while the spectral gap remains polynomially small. This comes from showing approximate tensorization of variance (Section 5.2.1).
Lemma 5.18 (Inverse-polynomial Spectral Gap at Criticality).
Let $G=(\mathcal{N},\mathcal{E})$ be a graph with constant maximum degree, $n$ vertices, and corresponding tree-unqiuness threshold for the $d$ and the hardcore model $\lambda_{c}$. The spectral gap of Glauber dynamics at fugacity $\lambda_{c}$ satisfies $\delta=\Omega(n^{-1-{4e\left(1+\frac{1}{d-2}\right)}})$.
Proof.
Follows from comments near the end of the [CCYZ25, Proof of Theorem 1.1]. ∎
In this case, we can apply Theorem 4.3 and obtain an asymptotically falling super quadratic speedup at criticality.
Theorem 5.19 (Bounded Degree Graphs At Criticality).
Let $P$ be a Glauber dynamics chain on a graph $G(\mathcal{N},\mathcal{E})$ with a bounded maximum degree $d=\mathcal{O}(1)$ at fugacity parameter $\lambda_{c}$ and stationary distribution $\pi_{\lambda_{c}}$. Also suppose that $\gamma$-spectral density is satisfied for $\gamma=\Theta(1)$. Then, there exists a short path algorithm that finds the maximum independent set in $G$ with running time
$\mathcal{O}^{\ast}\left([\pi_{\lambda_{c}}(E^{\ast})^{-1}]^{\frac{1}{2}-c(n)}\right),$
where $c(n)=\Omega(1/n^{1+4e\left(1+\frac{1}{d-2}\right)})$.
Proof.
Follows from Theorem 4.3, Lemma 5.18 and Lemma 5.15. ∎
As mentioned earlier, for random regular graphs, the Glauber dynamics mixes beyond $\lambda_{c}$ with high probability. This appears to hold all the way up to $\lambda=\mathcal{O}(1/\sqrt{d})$ [CCC+25]. Specifically, we have the following lemma extracted from [CCC+25].
Lemma 5.20 (Spectral Gap For Random Regular Beyond Criticality – Theorem 5.2 (2)[CCC+25]).
Let $G\sim\mathcal{G}_{\textup{RG}}(n,d)$ be a random $d$-regular graph for constant $d$. Consider the fugacity $\lambda=\frac{1-\xi}{2\sqrt{d-1}-1}$. Then with high probability $\pi_{\lambda}$ has a spectral gap lower bounded by $\delta=\Omega(\xi/n)$.
However, it appears that a lower bound was not provided fro the MLS or LS constants, i.e. [CCC+25, Theorem 5.2] becomes vacuous for the hardcore model. We again apply Theorem 4.3 and obtain a result like Theorem 5.19, providing an asymptotically falling super-quadratic speedup beyond the uniqueness threshold for random regular graphs.
Theorem 5.21 (Random Regular Graphs Beyond Criticality).
Let $P$ be a Glauber dynamics chain on a random $d$ regular-graph $G\sim\mathcal{G}_{\textup{RG}}(n,d)$, with constant $d$ and fugacity parameter $\lambda\leq\frac{1-\xi}{2\sqrt{d-1}-1}$, $\xi\in(0,1)$, and stationary distribution $\pi_{\lambda}$. Also suppose that $\gamma$-spectral density is satisfied for $\gamma=\Theta(1)$. Then with high probability over the choice of graph, there exists a short path algorithm that finds the maximum independent set in $G$ with running time
$\mathcal{O}^{\ast}\left([\pi_{\lambda}(E^{\ast})^{-1}]^{\frac{1}{2}-c(n)}\right),$
where $c(n)=\Omega(\xi n^{-1})$.
Proof.
Follows from Theorem 4.3, Lemma 5.20 and Lemma 5.15. ∎
Even with the above results, it is still not clear how Markov chain search compares to the best classical algorithms for MIS problem. As we consider a Markov chain that mixes to Gibbs distribution with fugacity $\lambda$, sampling from the Gibbs distribution may favor the smaller independent sets when $\lambda<1$. However, even if this is the case, the stationary distributions only has support on the constrained space (i.e. on the valid independent sets) rather than on the entire $2^{n}$ possible configurations. Therefore, we might still expect to prove that the success probability is larger than brute force search for certain graph models. To this end, the following proposition lower bounds the probability of finding the optimum by sampling from Glauber dynamics when restricting to random regular graphs.
@@
Let $\pi_{\lambda}$ be the stationary distribution of Glauber dynamics on a random $d$-regular graph $G\sim\mathcal{G}_{\textup{RG}}(n,d)$. Choose $x^{\ast}$ to be a particular maximum independent set in $G$. Then we have
$\pi_{\lambda}(x^{\ast})\geq 2^{-\kappa n},$
with $\kappa=-\frac{2\log(\lambda)\log d}{d}+\frac{1}{2d}+\frac{\log(1+\lambda)}{2}$.
@@
Note that for any constant $\xi\in(0,1)$, $\lambda=(1-\xi)\lambda_{c}$ and $\lambda_{c}=\frac{(d-1)^{(d-1)}}{(d-2)^{d}}$, we have that $\kappa<1$ for sufficiently large degree $d>d_{0}$ for some $d_{0}$. Hence the runtime is smaller than $2^{n}$, which is the runtime of brute force search over unconstrained space. Therefore, Markov chain search for this problem might be significantly faster than brute force search mainly due to fact that Glauber dynamics stays in the constrained space.
@@
The parameter $\kappa$ from Proposition 5.22 is a decreasing function of the degree. For sufficiently high degree, the runtime $2^{\kappa n}$ the is in fact faster than the best known generic algorithm for Maximum Independent Set (runtime of $\tilde{\mathcal{O}}(1.1996^{n})$, due to Xiao and Nagamuchi [XN17]).
5.2.3 Antiferromagnetic Ising Model
Consider the 2-spin Ising model on a graph $G=(\mathcal{N},\mathcal{E})$ defined via the Hamiltonian
$H(x)=\sum_{(i,j)\in\mathcal{E}}J_{ij}x_{i}x_{j}+\sum_{j}h_{j}x_{j},$
@@ 5.2.3 Antiferromagnetic Ising Model
We will consider the antiferromagnetic case $J_{ij}<0$ with no external-field. Specifically, let $J_{ij}=-1$ for all $(i,j)\in\mathcal{E}$ and $h=\mathbf{0}$. This model corresponds to Gibbs sampling with weights $\pi(x)\propto\exp(-\beta H(x))$, and a Gibbs sample can be prepared by using Glauber dynamics similar to the hardcore model. The only change is the transition probabilities, which can be computed by the marginal distribution $\pi(x_{i}^{t+1}|x^{t}_{\mathcal{N}\setminus\{i\}})$.
We consider the following short path Hamiltonian
@@
where $P_{\beta}$ is the Glauber dynamics transition matrix for inverse-temperature $\beta$. Similar to the setting of the MIS problem, a block-encoding of $D(P_{\beta})$ can be prepared efficiently.
First we have the following simple result bounding the psuedo-Lipschitz constant.
Lemma 5.24.
Let $P$ be a Glauber dynamics chain on the antiferromagnetic Ising model with inverse-temperature $\beta$ on a graph $G(\mathcal{N},\mathcal{E})$ with degree at most $d$. Then, $\|H\|_{P}=\mathcal{O}(d)$.
@@
$|H(x)-H(x^{\prime})|\leq 2d.$
Then, by Lemma 5.7, $\|H\|_{P}=\mathcal{O}(d)$. ∎
Corollary 5.25 (Log-Sobolev Constant at Constant-Fraction away From Criticality).
Let $G=(\mathcal{N},\mathcal{E})$ be any graph with constant-bounded degree $d$, $n$ vertices, and $\beta_{c}$, $\beta$ satisfies $(d-1)\tanh(\beta)\leq 1$ the corresponding tree-uniquness threshold for the antiferromagnetic Ising Model. If $\beta$ satisfies $(d-1)\tanh(\beta)\leq 1-\xi$ for any constant $\xi\in(0,1)$, then under the conditions of Theorem 5.12, the log-Sobolev constant of Glauber dynamics satisfies $\omega=\Omega(1/n)$.
Proof.
From [CLV21] and [CCYZ25, Lemma 3.24] for the specified $\beta$ range $\eta=\mathcal{O}(1)$, when $d$ and $\xi$ are constants. Hence Theorem 5.12 implies approximate entropy factorization with $C_{1}=\mathcal{O}(1)$. [CLV21, Proof of Theorem 1.1] gives that $\pi_{\beta}$ satisfies the bounded marginals condition when $d=\mathcal{O}(1)$. Hence, MLSI can be promoted to LSI when using approximate tensorization of entropy. Hence the LS constant is lower bounded by $\Omega(n^{-1})$. ∎
Theorem 5.26 (Constant Fraction away From Criticality).
Let $P$ be a Glauber dynamics chain for the antiferromagnetic Ising model over an arbitrary graph $G=(\mathcal{N},\mathcal{E})$ with constant maximum degree $d$. Suppose that for constant $\xi\in(0,1)$, the inverse temperature parameter $\beta$ satisfies $(d-1)\tanh(\beta)\leq 1-\xi$ and that the stationary distribution is $\pi_{\beta}$. If $\lvert E^{\ast}\rvert=\Theta(n)$, then there exists a short path algorithm that finds the ground state of the Ising model Hamiltonian running time bounded by
@@
We have that $\omega=\Omega(n^{-1})$ for specified $\beta$ range from Corollary 5.25. Furthermore, $\log(1/\pi_{\beta}(E^{\ast}))=\Theta(n)$ due to hardness of the problem. From Lemma 5.24 $\lVert H\rVert_{P}=\mathcal{O}(d)=\mathcal{O}(1)$. The rest follows by plugging in the parameters to Theorem 4.2. ∎
We also have the following result for at and slightly beyond criticality.
Corollary 5.27 (Log-Sobolev Constant at Criticality).
Let $G=(\mathcal{N},\mathcal{E})$ be any graph with constant-bounded degree $d$, $n$ vertices, and $\beta$ satisfies $(d-1)\tanh(\beta)\leq 1+o_{n}(1)$. Then under the conditions of Theorem 5.12, the log-Sobolev constant of Glauber dynamics satisfies $\omega=\Omega(1/n^{2+1/d})$.
Proof.
[CLV21, Proof of Theorem 1.1] gives that $\pi_{\beta}$ satisfies the bounded marginals condition when $d=\mathcal{O}(1)$. Hence, MLSI can be promoted to LSI when using approximate tensorization of entropy. From [CCYZ25, Proof of Theorem 3.23] we have that Glauber dynamics for the above specified range of $\beta$ satisfies approximate tensorization of entropy with $C_{1}=\mathcal{O}(n^{1+\frac{2}{d-2}})$. Hence the LS constant is lower bounded by $\Omega(n^{-2-1/d})$. ∎
Theorem 5.28 (Bounded Degree Graphs at Criticality).
Let $P$ be a Glauber dynamics chain for the antiferromagnetic Ising model over an arbitrary graph $G(\mathcal{N},\mathcal{E})$ with maximum constant degree $d$. Suppose the inverse temperature parameter $\beta$ satisfies $(d-1)\tanh(\beta)\leq 1+o_{n}(1)$ and that the stationary distribution is $\pi_{\beta}$. If $\lvert E^{\ast}\rvert=\Theta(n)$ and $\gamma$-spectral density is satisfied for $\gamma=\Theta(1)$, then there exists a short path algorithm that finds the ground state of the Ising model Hamiltonian running time bounded by
$\mathcal{O}^{\ast}\left([\pi_{\beta}(E^{\ast})^{-1}]^{\frac{1}{2}-c(n)}\right),$
where $c(n)=\Omega(n^{-1-1/d})$.
Proof.
From Corollary 5.27, we have a LS constant lower bounded by $\Omega(n^{-2-\frac{2}{d-2}})$. The rest of the proof follows the same as Theorem 5.26. ∎
We also have a super-quadratic speedup for the Ising model over random-regular graphs. However, unlike for the hardcore model, the speedup is an asymptotic super-quadratic speedup and does not fall with $n$. We start with a proposiiton bounding the size of $\lvert E^{\ast}\rvert$ over random regular graphs.
Proposition 5.29.
The optimum energy of Ising Model Hamiltonian $H$ on a random regular graph $G\sim\mathcal{G}(n,d)$ satisfies $|E^{\ast}|=\Theta(n)$ with high probability.
Proof.
Let $s$ denote the number of edges in the graph. The ground state of $H$ can be related to minimum bisection width [ZB10] denoted by $|\textrm{BW}|$ as follows
$|\textrm{BW}|=\frac{s+E_{\textrm{gs}}}{2}.$
Lemma 5.30 (Log-Sobolev Beyond Criticality For Random Regular).
Let $G\sim\mathcal{G}(n,d)$ be a random $d$-regular graph for constant $d$. Suppose the inverse temperature parameter $\beta$ satisfies $(d-1)\tanh(\beta)\leq\frac{d-1}{8\sqrt{d-1}-1}$. If $d=\mathcal{O}(1)$, then with high probability over the choice of graph, the log-Sobolev constant of Glauber dynamics for the antiferromagnetic Ising model satisfies $\omega=\Omega(1/n)$.
Proof.
[CLV21, Proof of Theorem 1.1] gives that $\pi_{\beta}$ satisfies the bounded marginals condition when $d=\mathcal{O}(1)$. Hence, MLSI can be promoted to LSI when using approximate tensorization of entropy. From [CCC+25, Theorem 5.2] we have MLS for $(d-1)\tanh(\beta_{c})\leq\frac{d-1}{8\sqrt{d-1}-1}$ scaling as $\Omega(\frac{e^{-\sqrt{d}}}{n})$ for constant $d$. ∎
Theorem 5.31 (Beyond Criticality for Random Regular Graphs).
Let $P$ be a Glauber dynamics chain for the antiferromagnetic Ising model over a randomly chosen $d$-regular graph $G\sim\mathcal{G}_{\textup{RG}}(n,d)$. Suppose the inverse temperature parameter $\beta$ satisfies $(d-1)\tanh(\beta)\leq\frac{d-1}{8\sqrt{d-1}-1}$ and that the stationary distribution is $\pi_{\beta}$. If $d=\mathcal{O}(1)$, then with high probability over the choice of graph there exists a short path algorithm that finds the ground state of the Ising model Hamiltonian running time bounded by
where $c>0$ is a constant.
Proof.
From Lemma 5.30, we have with high probability over the graph $\omega=\Omega(n^{-1})$. Furthermore, $\log(1/\pi_{\beta}(E^{\ast}))=\Theta(n)$ due to hardness of the problem. Again, $\lVert H\rVert_{P}=\mathcal{O}(d)=\mathcal{O}(1)$. Then the result follows from Corollary 4.5 applied to the case of Theorem 4.2. ∎
As mentioned earlier the inverse temperature bound, call it $\beta_{T}$, in Theorem 5.31 is very close to a known computational threshold, called the disorder chaos threshold. Specifically, it is known that with high probability, there exists a problem size $n_{0}$ such that no stable classical algorithm can mix in Wasserstein-2 distance when the inverse temperature exceeds $\frac{1}{\sqrt{d}}$ and $n\geq n_{0}$ [HPP24]. The quantum algorithm is improving upon a Gibbs sampler that mixes in TV to a point near this threshold. For a bounded domain, TV mixing implies Wasserstein-2 mixing. Unfortunately, the current provable edge that quantum has over the Gibbs sampler at the current known threshold $\beta_{T}<\frac{1}{\sqrt{d}}$ falls faster with $d$ than $\frac{1}{\sqrt{d}}-\beta_{T}$, implying it is currently unclear if quantum’s edge is enough to breach the Gibbs sampling threshold.
The following result rigorously shows that with high probability Markov chain search outperforms brute-force search for random regular graphs.
Proposition 5.32.
For the antiferromagnetic Ising model over random $d$-regular graphs, for any $\beta>0$, the probability of observing the ground state under the Gibbs measure is at least $2^{-\kappa n}$. With high probability in $n$, $\kappa>-1$.
Proof.
The replica symmetric bound for antiferromagnetic Ising over random regular graphs [COLMS22, Theorem 1.1] gives that
$\frac{\ln(Z(\beta))}{n}\leq\ln(2)+\frac{d}{2}\ln\left(\frac{1+e^{-\beta}}{2}\right)+\epsilon(n),$
with high probability in $n$. Hence
$\frac{\log(\pi(E^{\ast}))}{n}\geq\frac{\beta\lvert E^{\ast}\rvert}{n}-\frac{d}{2\ln(2)}\ln\left(\frac{1+e^{-\beta}}{2}\right)-1+\epsilon(n).$
If we apply Proposition 5.29, then the above is >1>-1 for $\beta>0,d\geq 0$ and sufficiently large $n$. ∎